OpenStudy (anonymous):

9x+4y=14 -6x+y=20 what is the solution method

myininaya (myininaya):

you try and i will check it

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

sorry

myininaya (myininaya):

its okay to be confused but show me an attempt and how i will tell you where u are going wrong

OpenStudy (anonymous):

multiply the second part by -4 9x+4y=14 24x-4y=-80 then 9x+4y=14 + 24x-4y=-80 =32x=-64 =x=-64/32 x=-2 9(-2)+4y=14 -18+4y=14 4y=14+18 y =32/4 y=8 the answer is x=-2 y=8

OpenStudy (anonymous):

The process of above answer is wrong.Because if you want to multiply with -4 you need to do so in both the equations.better to solve the sum through substitution method. y=20+6x........(1) 9x+4(20+6x)=14 or,9x+80+24x=14 or,33x=-66 x=-2 y=20-12=8